What is the osmotic pressure of 0.9% m/v solution of NaCl at 25°C?


1
Expert's answer
2021-05-10T01:51:05-0400

Osmotic pressure "= i \times C \times R \times T"


where, C= concentration of solute(in terms of Molarity)


"R = 0.082"


"i = 1"


0.9% "NaCl" solution means 0.9g is present in 100mL of solution.


Moles of "NaCl" "= \dfrac{0.9}{58} = 0.015"


Concentration of "NaCl = \dfrac{0.015}{100} \times 1000 = 0.10"


Hence, Osmotic pressure "= 1 \times 0.10 \times 0.082 \times 300"


"= 2.46atm"